# The Very Faint X-ray Transient 4XMM J174610.7–290020 at the Galactic center

### v1

@@ 2.3 Chandra
There are no Chandra (Weisskopf et al., 2000) observations that cover the outburst in 2024. However, the same source, 4XMM J174610.7–290020, is also reported in the Chandra Source Catalog (CSC 2.1, Evans et al., 2024) with identifier 2CXO J174610.7–290019; as CXOU J174610.8–290019 in the Chandra catalog of the GC (Muno et al., 2009) and in the Ultradeep Chandra Catalog (Zhu et al., 2018, source n. 3596). According to the CSC 2.1, there are 11 observations between 1999 and 2024 in which the source has been detected with a significance ¿ 3444The significance is the ratio of the photon flux measurement to its average error, see cxc.cfa.harvard.edu/csc/columns/significance.html for details. in the ACIS-I 2.0–7.0 keV band.
@@ 2.4 NuSTAR
NuSTAR (Harrison et al., 2013) detected the source during the late stages of the 2024 outburst. However, given the relatively low flux, the NuSTAR analysis does not add a significant contribution to the XMM-Newton observations of the outburst. On the other hand, the 2025 outburst was detected in four NuSTAR observations. We collected the spectra of the source in a circular region of radius 20″, centered on the source. The background is accumulated in a 70 arcsec circular region at the coordinates: RA, Dec: (17h46m07.8s, -2901’53.3”). When analyzing the NuSTAR spectra we fit the two modules FPMA and FPMB simultaneously, including a global multiplicative constant to account for cross-calibration uncertainties between the two instruments. We verify that the two spectra are overall consistent within a factor of 10% in all observations.
@@ 2.5 Einstein Probe (EP)
EP detected the outburst of this source in late March 2024 during a performance verification (PV) observation of the GC, using the Follow-up X-ray Telescope (FXT) with an exposure time of $\sim$ 31.6 ks. EP-FXT comprises two co-aligned modules (FXT-A and FXT-B) in the 0.3-10 keV band, and the observation data of these two modules are processed using the FXT data analysis software (fxtsoftware v1.10).
@@ 3.1 Light curves
Figure 2: Light curve of the outburst during 2024 (left) and 2025 (right). $F_{2-10}$ is the 2–10 keV flux corrected for absorption, and it is computed assuming an absorbed power law spectrum, with $N_{\mathrm{H}}=2\times 10^{23}$ cm-2 and $\Gamma=2$. The two XRISM data are compute from the count rates reported in Yoshimoto et al. (2025), assuming the same spectral model.
Figure 3: XMM-Newton EPIC-pn light curves of the 2024 outburst (blue data), in the 2–10 keV band. Below each light curve, red data points are the hardness ratio computed from the light curves in the 2–5 and 5–10 keV energy bands. The gray dashed line is the median hardness ratio for each observation.
The left panel of Fig. 2 displays the light curve obtained during the outburst of the source in 2024. The count rate of the various instruments is converted to a 2–10 keV flux using webpimms [^5] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl, assuming a spectral model of an absorbed, powerlaw with a photon index $\Gamma=2$, and $N_{\mathrm{H}}=2\times 10^{23}$ cm-2 (see, Sect. 3.2). The duration of the outburst is $\sim 50$ days. A gap in the Swift-XRT light curve is due to the fact that the satellite was in safe-mode from 15-03-2024 to 03-04-2024, for an issue related to the gyroscopes.
@@
Table 3: Timing analysis of the five XMM-Newton observations. The second column lists the $\chi^{2}$ relative to a fit using the weighted mean as a constant; in parenthesis is the number of data points.
The 2-10 keV light curves of the five XMM-Newton observations performed in spring 2024 are shown in Fig. 3 (blue), binned to 800 s. The last light curve, taken roughly six days after the others, exhibits a count rate reduced by a factor of about 4 compared to the previous four observations. All five light curves show some variability. To test whether the variability is statistically significant, we calculated the $\chi^{2}$ with respect to the weighted mean of each observation. The results are shown in Table 3. In the same table, we report the ratio of the standard deviation and the weighted mean. Overall, the flux is fluctuating within 25-50% of its mean value.
The hardness ratio [^6] The hardness ratio (HR) is defined by (HS)/(H+S)(H-S)/(H+S), where $H$ and $S$ are the count rates in the hard (5–10 keV) and soft (2–5 keV) bands, respectively. between the 5-10 and 2-5 keV light curves shown in Fig. 3 does not provide significant evidence for spectral variations.
We looked for periodicity of the emission in the 2–10 keV band by computing the Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982) in the frequency range $4\times 10^{-5}-10^{-1}$ Hz. The periodogram of the longest XMM-Newton observation (ID. 0944580701, 2024-03-27) is shown as an example in Fig. 4; the other periodograms have similar shapes. Yoshimoto et al. (2025) found a potential hint of periodicity of $\simeq 1537$ s, which we do not observe in the XMM-Newton data.
Figure 4: Lomb-Scargle periodogram of XMM-Newton obs. 0944580701 (2024-03-27). The other observations show a similar behavior.
@@ 3.2 Spectral analysis
Figure 5: XMM-Newton EPIC-pn spectrum of the source (blue) accumulated in Obs. 0944580701, with the 3 models discussed in Sect. 3.2. The spectra accumulated in the other 4 XMM-Newton observations present similar shape. Model 1 is an absorbed power law with the Fe XXV (6.7 keV) and Fe XXVI (6.97 keV) lines, while model 2 includes a single reflection line. The lines are highlighted with dotted lines. Model 3 is ionized thermal plasma component (apec). The dashed orange line is the modeled background contribution.
@@
Figure 5 displays the EPIC-pn spectrum of the longest XMM-Newton observation (2024-03-27). The spectra of the other four XMM-Newton observations, and the EP observations, present a similar shape. The spectrum is characterized by an excess at $\sim$ 7 keV. The most prominent emission lines in this region are the highly ionized iron lines Fe XXV ($\simeq 6.7$ keV) and Fe XXVI ($\simeq 6.97$ keV).
@@
Since the five XMM-Newton light curves exhibit different fluxes, and to maintain full generality, we fit each observation independently. As a first model (”model 1”), we implemented an X-ray-halo-scattered, absorbed power law with two Gaussian lines corresponding to the Fe XXV and Fe XXVI lines (fgcdust*tbabs*(powerlaw+gaussian+gaussian) in Xspec). The X-ray halo scattered model (fgcdust) is described in Jin et al. (2017, 2018). This model is adopted since the source, as we will show, is heavily absorbed, and in the projected direction of the GC, therefore it is likely a few kpc from the observer. This particular dust scattering model accounts for the dust material in the spiral arms of the Milky Way. We assumed abundances from Wilms et al. (2000), and cross-section from Verner et al. (1996). The best parameters, along with their uncertainties, are reported in Table 4, while the spectrum is displayed in the left panel of Fig. 5. The source appears heavily absorbed with a column density typical of the sources located at the GC ($N_{\mathrm{H}}\sim 10^{23}$ cm-2, Baganoff et al., 2003). The photon index is consistent with $\Gamma\simeq 2$. The Fe XXVI line is the most prominent of the two iron lines and is consistently detected; while on two occasions we where able to only measure an upper limit for the flux of the Fe XXV line (see Table 4).
@@
As an alternative model, since the Fe XXV and XXVI lines are not resolved (see Fig. 5), we substitute the two ionized lines with a single line, allowing a non-null width (”model 2”, second panel in Fig. 5). This is the case of a single reflection line from the accretion disk (Fabian et al., 1989). Table 4 reports the fit result. Both the $N_{\mathrm{H}}$ and $\Gamma$ parameters are not affected when switching from a model with to lines to that with a single one. The energy of the line is found to be 6.86.96.8-6.9 keV, whereas the fit returns an upper limit on the width of $\simeq 200-300$ eV. The flux of the line is overall consistent among the first four observations, with a slight decrease in the last one. In the same table, we report the difference of the Akaike information criterion (AIC) between model 2 and model 1. The model with the two ionized lines is slightly preferred, but the single reflection line cannot be ruled out.
@@
Given the previous results, we attempt a fit of the spectrum with a collisionally-ionized diffuse gas [^7] A description of the apec model is available at http://atomdb.org/ (fgcdust*tbabs*apec in Xspec, ”model 3”). Details of this fit are displayed in Table 5, and the spectrum is shown in the right panel of Fig. 5. The measured absorption column density is lower than the one inferred in the previous fits. The estimated temperature, which is set by the energy of the ionized iron lines and the continuum shape, is $kT\sim 12$ keV. In the table, we report the difference between the AIC of model 3 and model 1. Apart from the observation taken on 2024-04-04, there is no clear preference between the two models.
@@
Table 4: Best fit parameters and associated uncertainties ($1\,\sigma$) for the five XMM-Newton spectra extracted during the 2024 outburst. Model 1 includes the Fe XXV and Fe XXVI lines, while model 2 a single reflection line. $F_{2-10}$ is the de-absorbed flux in the 2-10 keV band, assuming model 1.
[…]
@@
Table 5: Same as Tab. 4 but with an apec model instead of a power law model.
4 Rebrightening of 2025
On 04 April 2025, NuSTAR serendipitously detected a rebrightening of the source (Grefenstette et al., 2025). The new outburst was confirmed the following day by Swift-XRT. The light curve of this second outburst is displayed in the right panel of Fig. 2. The source is clearly detected between 2025-04-04 and 2025-04-09 in four NuSTAR and three Swift-XRT observations. The source was not detected by Swift-XRT on 2025-04-02, and in the days after 2025-04-10. Therefore, the duration of this second outburst is significantly shorter than the 2024 event, elapsing only $\simeq 5$ days.
We collected the spectrum during the four NuSTAR observations in which the transient was detected, and performed the fit in the 3-20 keV band. Given that the NuSTAR spectral resolution is lower than XMM-Newton, we fit only the model 2 and model 3, without attempting to distinguish between a single or a combination of two iron lines. Figure 6 shows the NuSTAR FPMB spectrum taken on 2025-04-07 (ID. 31002004006), along with the two models. The best fit parameters are reported in Table 6. To allow a direct comparison between the parameters of the 2024 outburst, and since NuSTAR is not sensitive below 3 keV, we fixed the column density to the best values found in the previous outburst. Overall, the spectrum appears softer when compared with the 2024 outburst (Table 4 and 5). If fitted with a power law, the photon index is close to $\Gamma\simeq 3$, while fitted with an apec component, the temperature is $kT\simeq 5-6$ keV. The normalization of the Fe line is consistent between the two outbursts. The AIC difference clearly shows a preference for model 2 (powerlaw + line). The apec model, on the other hand, struggles to reproduce simultaneously the iron lines and the continuum.
@@ 4 Rebrightening of 2025
Table 6: Best spectral fit parameters of the 2025 outburst, observed with NuSTAR.
: fixed.
Figure 6: NuSTAR FPMB spectrum of the source (blue) accumulated in Obs. 31002004006, with the 2 models discussed in Sect. 4. Model 2 includes a single reflection line, with a power law continuum. The line is highlighted with a dotted line. Model 3 is ionized thermal plasma component (apec). The dashed orange line is the background contribution.
@@ 5 Past outbursts and detections
In Fig. 7, we present the light curve of the source from 2000 onward, compiled using archived observations from Chandra. Upper limits (blue data points) and detections (orange) are retrieved from the CSC 2.1. The conversion from count rate to a flux is performed with the same model used in Fig. 2 ($N_{\mathrm{H}}=2\times 10^{23}$ cm-2 and $\Gamma=2$). For graphical purposes, consecutive observations that are within 120 days are merged. The mean flux is shown in cases of detection, while only the more stringent upper limit is shown in cases of non-detection. The source has been detected by Chandra a few times, with a 2-10 keV flux within $2\times 10^{-14}$ erg s-1 cm-2 and $4\times 10^{-13}$ erg s-1 cm-2. The reported upper limits indicate a quiescence flux $F_{2-10}\leq 4\times 10^{-15}$ erg s-1 cm-2.
@@ 5.1 Potential type I burst
During an XMM-Newton observation performed on 2004-08-31 (ID n. 0202670701), in which the source was detected along the whole observations, the light curve of the source displayed a potential type I burst (Pastor-Marazuela et al., 2020). Figure 8 displays the histogram of the events (bin-size of 10 s) during the burst. To determine the duration of the burst, we performed a Bayesian-block decomposition (Scargle et al., 2013) on the EPIC pn unfiltered event list associated with the source. We used the bayesian_blocks function from the Python package Astropy, assuming $p0=0.01$, where $p0$ gives the false alarm probability to compute the prior. The burst is clearly identified for a total duration of 523 s. The duration is longer than typical X-ray type I bursts (from a few seconds to about a minute, Galloway et al., 2008), but consistent with “intermediate-duration bursts” that last a few to tens of minutes (Alizai et al., 2023). The peak of the burst is reached after $\simeq 40$ s.
We integrated the spectrum during the flare interval in the same region as before. We estimated the background by accumulating a spectrum in the same position, before the start of the burst feature. We chose a phenomenological model for the background (instrumental and pre-burst emission) consisting of an absorbed thermal plasma component, with a power law and a 6.4 keV line (tbabs*(apec+pow+gauss) in Xspec). The model for the burst is an absorbed black body emission (fgcdust*tbabs*bbodyrad in Xspec). The spectrum, with the associated fit, is shown in Fig. 9. The derived temperature is $kT=2.8^{+1.4}_{-0.7}$ keV, for an absorption column density of $N_{\mathrm{H}}=(7\pm 5)\times 10^{22}$ cm-2. The mean deabsorbed flux in the 2-10 keV band is $F_{2-10}=1.9^{+0.3}_{-0.2}\times 10^{-11}$ erg s-1 cm-2. For a given distance $d$, the normalization of the black body spectrum can be converted to the source radius ($R$) by:
@@
In order to search for any spectral softening during the burst, we divided it into three intervals highlighted in Fig. 8. We then fitted the spectrum in each interval; the best parameters are reported in Table 7. A softening of the spectrum is identified, consistent with the typical behavior of type I X-ray bursts. To verify that the spectrum is not excessively affected by the pile-up effect during the $t_{1}$ interval, we performed the analysis again, selecting only single-pattern events (”PATTERN==0” in the SAS evselect function). The results are consistent within the statistical uncertainties.
Type I bursts which reach Eddington luminosity can be exploited as standard candles to determine the distance to the NS-LMXB. This kind of type I bursts present photospheric radius expansion, and result in a bolometric peak luminosity of $L_{\mathrm{bol}}\simeq 3.8\times 10^{38}$ erg s-1 (Kuulkers et al., 2003). The burst of 2004, that has no evidence of photospheric radius expansion, is much fainter. The peak lies within the $t_{1}$ interval, as shown in Fig. 8 and the count rate of the peak is about twice the mean count rate measured during $t_{1}$. Assuming that the source is located at the GC (distance of 8.2 kpc), the luminosity in the 2-10 keV band during $t_{1}$ is $L_{2-10}=(7\pm 1)\times 10^{35}$ erg s-1, which translates into a bolometric luminosity peak luminosity of $L_{\mathrm{bol}}\sim 10^{37}$ erg s-1, computed in the energy range 0.1-100 keV.
@@
Figure 9: XMM-Newton EPIC-pn spectrum of the potential type I burst occurred in 2004. The model is an absorbed black body (see Table 7).
@@
Table 7: Best fit parameters for the potential type I burst.
: fixed.
[…]

### v2

@@ 2.3 Chandra
There are no Chandra (Weisskopf2000) observations that cover the outburst in 2024. However, the same source is also reported in the Chandra Source Catalog (CSC 2.1, Evans2024) with identifier 2CXO J174610.7-290019; as CXOU J174610.8-290019 in the Chandra catalog of the GC (Muno2009) and in the Ultradeep Chandra Catalog (Zhu2018, source n. 3596). According to the CSC 2.1, there are 11 observations between 1999 and 2024 in which the source has been detected with a significance ¿ 3666The significance is the ratio of the photon flux measurement to its average error; see cxc.cfa.harvard.edu/csc/columns/significance.html for details. in the ACIS-I 2.0–7.0 keV band.
@@ 2.4 NuSTAR
NuSTAR (Harrison2013) detected the source during the late stages of the 2024 outburst. However, given the relatively low flux, the NuSTAR analysis does not add a significant contribution to the XMM-Newton observations of the outburst. On the other hand, the 2025 outburst was detected in four NuSTAR observations. We collected the spectra of the source in a circular region of radius 20 arcsec, centered on the source. The background is accumulated in a 70 arcsec circular region at the coordinates: RA, Dec: (17h46m07.8s, -2901’53.3”). When analyzing the NuSTAR spectra, we fit the two modules FPMA and FPMB simultaneously, including a global multiplicative constant to account for cross-calibration uncertainties between the two instruments. We verified that the two spectra are overall consistent within a factor of 10% in all observations.
@@ 2.5 Einstein Probe (EP)
The Einstein Probe detected the outburst of this source in late March 2024 during a performance verification observation of the GC, using the Follow-up X-ray Telescope (FXT) with an exposure time of $\sim$ 31.6 ks. EP-FXT comprises two co-aligned modules (FXT-A and FXT-B) in the 0.3-10 keV band, and the observation data of these two modules are processed using the FXT data analysis software (fxtsoftware v1.10).
@@ 3.1 Light curves
Figure 2: Light curve of the outburst in 2024 (left) and 2025 (right). $F_{2-10}$ is the 2–10 keV flux corrected for absorption, and it is computed assuming an absorbed power law spectrum, with $N_{\mathrm{H}}=2\times 10^{23}$ cm-2 and $\Gamma=2$. The two XRISM data points were computed from the count rates reported in Yoshimoto2025, assuming the same spectral model.
Figure 3: XMM-Newton EPIC-pn light curves of the 2024 outburst (blue data) in the 2–10 keV band. Below each light curve, red data points are the hardness ratio computed from the light curves in the 2–5 and 5–10 keV energy bands. The dashed gray line is the median hardness ratio for each observation.
The left panel of Fig. 2 displays the light curve obtained during the outburst of the source in 2024. The count rate of the various instruments is converted to a 2–10 keV flux using webpimms [^7] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl, assuming a spectral model of an absorbed, power law with a photon index $\Gamma=2$, and $N_{\mathrm{H}}=2\times 10^{23}$ cm-2 (see Sect. 3.2). The duration of the outburst is $\sim 50$ days. A gap in the Swift-XRT light curve is due to the fact that the satellite was in safe-mode from 15-03-2024 to 03-04-2024, for an issue related to the gyroscopes.
Table 3: Timing analysis of the five XMM-Newton observations. [^8] The second column lists the $\chi^{2}$ relative to a fit using the weighted mean as a constant; in parenthesis is the number of data points.
@@
The 2–10 keV light curves of the five XMM-Newton observations performed in spring 2024 are shown in Fig. 3 (blue), binned to 800 s. The last light curve, taken roughly six days after the others, exhibits a count rate reduced by a factor of about 4 compared to the previous four observations. All five light curves show some variability. To test whether the variability is statistically significant, we calculated the $\chi^{2}$ with respect to the weighted mean of each observation. The results are shown in Table 8. In the same table, we report the ratio of the standard deviation and the weighted mean. Overall, the flux is fluctuating within 25-50% of its mean value. The hardness ratio [^9] The hardness ratio (HR) is defined by (HS)/(H+S)(H-S)/(H+S), where $H$ and $S$ are the count rates in the hard (5–10 keV) and soft (2–5 keV) bands, respectively. between the 5–10 and 2–5 keV light curves shown in Fig. 3 does not provide significant evidence of spectral variations.
We looked for periodicity of the emission in the 2–10 keV band by computing the Lomb-Scargle periodogram (Lomb1976; Scargle1982) in the frequency range $4\times 10^{-5}-10^{-1}$ Hz. The periodogram of the longest XMM-Newton observation (ID. 0944580701, 2024-03-27) is shown as an example in Fig. 4; the other periodograms have similar shapes. Yoshimoto2025 found a potential hint of periodicity of $\simeq 1537$ s, which we do not observe in the XMM-Newton data.
Figure 4: Lomb-Scargle periodogram of XMM-Newton obs. 0944580701 (27 March 2024). The other observations show a similar behavior.
@@ 3.2 Spectral analysis
Figure 5: XMM-Newton EPIC-pn spectrum of the source (blue) accumulated in Obs. 0944580701, with the three models discussed in Sect. 3.2. The spectra accumulated in the other four XMM-Newton observations present similar shapes. Model 1 is an absorbed power law with the Fe XXV (6.7 keV) and Fe XXVI (6.97 keV) lines (dotted lines), while model 2 includes a single reflection line. Model 3 is ionized thermal plasma component (apec). The dashed orange line is the modeled background contribution.
@@
Figure 5 displays the EPIC-pn spectrum of the longest XMM-Newton observation (27 March 2024). The spectra of the other four XMM-Newton observations, and the EP observations, present a similar shape. The spectrum is characterized by an excess at $\sim$ 7 keV. The most prominent emission lines in this region are the highly ionized iron lines Fe XXV ($\simeq 6.7$ keV) and Fe XXVI ($\simeq 6.97$ keV).
@@
Since the five XMM-Newton light curves exhibit different fluxes, and to maintain full generality, we fit each observation independently. As a first model (”model 1”), we implemented an X-ray-halo-scattered, absorbed power law with two Gaussian lines corresponding to the Fe XXV and Fe XXVI lines (fgcdust*tbabs*(powerlaw+gaussian+gaussian) in Xspec). The X-ray halo scattered model (fgcdust) is described in Jin2017; Jin2018. This model is adopted since the source, as we will show, is heavily absorbed and in the projected direction of the GC; it is therefore likely a few kiloparsecs from the observer. This particular dust scattering model accounts for the dust material in the spiral arms of the Milky Way. We assumed abundances from Wilm2000, and cross section from Verner1996. The best parameters, along with their uncertainties, are reported in Table 11, while the spectrum is displayed in the left panel of Fig. 5. The source appears heavily absorbed with a column density typical of the sources located at the GC ($N_{\mathrm{H}}\sim 10^{23}$ cm-2, Baganoff2003). The photon index is consistent with $\Gamma\simeq 2$. The Fe XXVI line is the most prominent of the two iron lines and is consistently detected; while on two occasions we where able to only measure an upper limit for the flux of the Fe XXV line (see Table 11).
@@
As an alternative model, since the Fe XXV and XXVI lines are not resolved (see Fig. 5), we substituted the two ionized lines with a single line, allowing a non-null width (”model 2”; second panel in Fig. 5). This is the case of a single reflection line from the accretion disk (Fabian1989). Table 11 reports the fit result. Neither the $N_{\mathrm{H}}$ nor $\Gamma$ parameters were affected when switching from a model with two lines to that with a single one. The energy of the line is found to be 6.86.96.8-6.9 keV, whereas the fit returns an upper limit on the width of $\simeq 200-300$ eV. The flux of the line is overall consistent among the first four observations, with a slight decrease in the last one. In the same table, we report the difference of the Akaike information criterion (AIC) between model 2 and model 1. The model with the two ionized lines is slightly preferred, but the single reflection line cannot be ruled out.
@@
Given the previous results, we attempted a fit of the spectrum with a collisionally ionized diffuse gas [^10] A description of the apec model is available at http://atomdb.org/ (fgcdust*tbabs*apec in Xspec; ”model 3”). Details of this fit are displayed in Table 5, and the spectrum is shown in the right panel of Fig. 5. The measured absorption column density is lower than the one inferred in the previous fits. The estimated temperature, which is set by the energy of the ionized iron lines and the continuum shape, is $kT\sim 12$ keV. In the table, we report the difference between the AIC of model 3 and model 1. Apart from the observation taken on 4 April 2024, there is no clear preference between the two models.
@@
Table 4: Best-fit parameters and associated uncertainties ($1\,\sigma$) for the five XMM-Newton spectra extracted during the 2024 outburst. [^11] Model 1 includes the Fe XXV and Fe XXVI lines, while model 2 a single reflection line. $F_{2-10}$ is the de-absorbed flux in the 2-10 keV band, assuming model 1.
[…]
@@
Table 5: Same as Table 11 but with an apec model instead of a power law model.
@@
4 Re-brightening of 2025
On 4 April 2025, NuSTAR serendipitously detected a re-brightening of the source (Grefenstette2025). The new outburst was confirmed the following day by Swift-XRT. The light curve of this second outburst is displayed in the right panel of Fig. 2. The source is clearly detected between 4 and 9 April 2025 in four NuSTAR and three Swift-XRT observations. The source was not detected by Swift-XRT on 2 April 2025, and in the days after 10 April 2025. Therefore, the duration of this second outburst is significantly shorter than the 2024 event, elapsing only $\simeq 5$ days.
We collected the spectrum during the four NuSTAR observations in which the transient was detected, and performed the fit in the 3-20 keV band. Given that the NuSTAR spectral resolution is lower than XMM-Newton, we fit only the model 2 and model 3, without attempting to distinguish between a single or a combination of two iron lines. Figure 6 shows the NuSTAR FPMB spectrum taken on 7 April 2025 (ID. 31002004006), along with the two models. The best-fit parameters are reported in Table 12. To allow a direct comparison between the parameters of the 2024 outburst, and since NuSTAR is not sensitive below 3 keV, we fixed the column density to the best values found in the previous outburst. Overall, the spectrum appears softer when compared with the 2024 outburst (Table 11 and 5). If fitted with a power law, the photon index is close to $\Gamma\simeq 3$, while fitted with an apec component, the temperature is $kT\simeq 5-6$ keV. The normalization of the Fe line is consistent between the two outbursts. The AIC difference clearly shows a preference for model 2 (power law + line). The apec model, on the other hand, struggles to reproduce simultaneously the iron lines and the continuum.
Table 6: Best spectral fit parameters of the 2025 outburst, observed with NuSTAR.121212 fixed.
@@ 4 Re-brightening of 2025
Figure 6: NuSTAR FPMB spectrum of the source (blue) accumulated in Obs. 31002004006, with the two models discussed in Sect. 4. Model 2 includes a single reflection line (dotted line) with a power law continuum. Model 3 is an ionized thermal plasma component (apec). The dashed orange line is the background contribution.
@@ 5 Past outbursts and detections
In Fig. 7 we present the light curve of the source from 2000 onward, compiled using archived observations from Chandra. Upper limits (blue data points) and detections (orange) are retrieved from the CSC 2.1. The conversion from count rate to a flux is performed with the same model used in Fig. 2 ($N_{\mathrm{H}}=2\times 10^{23}$ cm-2 and $\Gamma=2$). For graphical purposes, consecutive observations that are within 120 days are merged. The mean flux is shown in cases of detection, while only the more stringent upper limit is shown in cases of non-detection. The source has been detected by Chandra a few times, with a 2–10 keV flux within $2\times 10^{-14}$ erg s-1 cm-2 and $4\times 10^{-13}$ erg s-1 cm-2. The reported upper limits indicate a quiescence flux $F_{2-10}\leq 3\times 10^{-15}$ erg s-1 cm-2.
@@ 5.1 Potential type I burst
During an XMM-Newton observation performed on 31 August 2004 (ID n. 0202670701), in which the source was detected along the whole observations, the light curve of the source displayed a potential type I burst (Pastor-Marazuela2020). Figure 8 displays the histogram of the events (bin-size of 10 s) during the burst. To determine the duration of the burst, we performed a Bayesian-block decomposition (Scargle2013) on the EPIC pn unfiltered event list associated with the source. We used the bayesian_blocks function from the Python package Astropy, assuming $p0=0.01$, where $p0$ gives the false alarm probability, which is used to compute the prior. The burst is clearly identified for a total duration of 523 s. The duration is longer than typical X-ray type I bursts (from a few seconds to about a minute, Galloway2008), but consistent with “intermediate-duration bursts” that last a few to tens of minutes (Alizai2023). The peak of the burst is reached after $\simeq 40$ s.
We integrated the spectrum during the flare interval in the same region as before. We estimated the background by accumulating a spectrum in the same position, before the start of the burst feature. We chose a phenomenological model for the background (instrumental and pre-burst emission) consisting of an absorbed thermal plasma component, with a power law and a 6.4 keV line (tbabs*(apec+pow+gauss) in Xspec). The model for the burst is an absorbed black body emission (fgcdust*tbabs*bbodyrad in Xspec). The spectrum, with the associated fit, is shown in Fig. 9. The derived temperature is $kT=2.8^{+1.4}_{-0.7}$ keV, for an absorption column density of $N_{\mathrm{H}}=(7\pm 5)\times 10^{22}$ cm-2. The mean de-absorbed flux in the 2-10 keV band is $F_{2-10}=1.9^{+0.3}_{-0.2}\times 10^{-11}$ erg s-1 cm-2. For a given distance ($d$), the normalization of the black body spectrum can be converted to the source radius ($R$) as
@@
To search for any spectral softening during the burst, we divided it into three intervals highlighted in Fig. 8. We then fit the spectrum in each interval; the best parameters are reported in Table 13. A softening of the spectrum is identified, consistent with the typical behavior of type I X-ray bursts. To verify that the spectrum is not excessively affected by the pile-up effect during the $t_{1}$ interval, we performed the analysis again, selecting only single-pattern events (”PATTERN==0” in the SAS evselect function). The results are consistent within the statistical uncertainties.
Type I bursts that reach Eddington luminosity can be exploited as standard candles to determine the distance to the NS-LMXB. This kind of type I burst presents photospheric radius expansion and has a bolometric peak luminosity ($L_{\mathrm{bol}}$) of $\simeq 3.8\times 10^{38}$ erg s-1 (Kuulkers2003). The burst of 2004, for which there is no evidence of photospheric radius expansion, is much fainter. The peak lies within the $t_{1}$ interval, as shown in Fig. 8, and the count rate of the peak is about twice the mean count rate measured during $t_{1}$. Assuming that the source is located at the GC (distance of 8.2 kpc), the luminosity in the 2–10 keV band during $t_{1}$ is $L_{2-10}=(7\pm 1)\times 10^{35}$ erg s-1, which translates into a bolometric peak luminosity of $L_{\mathrm{bol}}\sim 10^{37}$ erg s-1, computed in the energy range 0.1–100 keV.
@@
Figure 9: XMM-Newton EPIC-pn spectrum of the potential type I burst that occurred in 2004. The model is an absorbed black body (see Table 13).
Table 7: Best-fit parameters for the potential type I burst.131313: fixed.
[…]